Find the height of a trapezoid with area of 99m2 and the bottom base is twice the top base, which measures 6m.


1
Expert's answer
2021-12-16T17:25:53-0500

SOLUTION


Area of trapezoid"=\frac{a+b}{2}h"

If the bottom base is twice the top base then,

The top base is "6m"

The bottom base is "6\times2=12m"

Thus,

"99m^2=\frac{12+6}{2}h"

"99m^2=\frac{18}{2}h"

"99=9h"

"h=\frac{99}{9}"

"h=11m"

Therefore, the height of the trapezoid is 11m.